# Infinite-Width Limit of a Single Attention Layer: Analysis via Tensor Programs

### v2

@@ 5.2 Scalings of the Dot-product Score and Finiteness of Heads
Figure 2: Visualization of the dot-product score $p_{1,1}^{(1)}$ and attention output $y_{1}^{1}$, as defined in Example 3.1, comparing finite-width behavior to their infinite-width limits. (a) Histogram of the empirical distribution of $p_{1,1}^{(1)}$ for $n=256$ alongside the plot of its infinite-width limit distribution $\mathring{p}_{1,1}^{1}$, under two scaling schemes; $1/\sqrt{n}$ and $1/n$. The $1/n$-scaled score collapses to zero in the infinite-width limit, while the $1/\sqrt{n}$-scaled score retains a nondegenerate distribution. (b) Kernel density estimates of the empirical distribution of $y_{1}^{1}$ for $n=256$ (dashed lines) alongside the plot of its infinite-width limit distribution $Z^{y^{1}}$ (solid lines), varying head counts $H\in\{1,256\}$. The black solid line represents the density of the infinite-head limit distribution from HBSDN (20). This demonstrates that our theoretical prediction remains accurate even when $H$ grows, and it approaches the infinite-head limit.
We next investigate two facets of finite-width behavior in Example 3.1, which are the effect of different scaling rules on the variability of the dot-product score and the impact of a finite number of heads on the attention output. For comparison, Yan19b assumes $1/n$-scaling, and the infinite-head result by HBSDN (20) applies only in the limit $n,H\to\infty$.
@@
Figure 2(b) reports histograms of the attention output $y_{1}^{1}$ for $n=256$ with head counts $H=1$ and $H=256$, overlaid with the infinite-width densities of $Z^{y^{1}}$, which are approximated via Monte Carlo sampling. The close agreement between finite-width histograms and their infinite-width limit curves, even when $H$ grows on the same order as $n$, demonstrates the robustness of our infinite-width approximation in the growing number of heads. Furthermore, it is observed that as $H$ increases, our non-Gaussian distribution approaches the Gaussian distribution from HBSDN (20).
@@ 6 Conclusion
Looking forward, we believe our framework can serve as a foundation for future extension to deep Transformer architectures. We predict that as layers are stacked, the output distribution may increasingly diverge from any hierarchical Gaussian form, possibly yielding even heavier-tailed non-Gaussian behaviors. This departure from Gaussianity is not merely a theoretical curiosity, but it has profound implications for the learning dynamics of attention-based models. For instance, the presence of higher-order moments associated with such non-Gaussianity could influence signal propagation by preserving feature variability across layers, thereby reducing the risk of rank collapse. Furthermore, the anisotropy induced by non-Gaussianity may lead to more irregular curvature in the optimization landscape, possibly affecting the convergence properties of training dynamics (e.g., through sharper gradients or more prominent saddle regions). From a feature learning perspective, heavier-tailed distributions might even prove advantageous, as amplified high-magnitude components could facilitate the emergence of dominant features. Consequently, this prediction, that the divergence from Gaussianity intensifies as layers are stacked, suggests that a new framework distinct from existing Gaussian-based approaches is essential for analyzing deep architectures. Our work provides an important first step towards this direction.
@@ Acknowledgements
Mana Sakai was supported by RIKEN Junior Research Associate Program. Ryo Karakida was supported by JST FOREST (Grant No. JPMJFR226Q) and JSPS KAKENHI (Grant No. 22H05116). Masaaki Imaizumi was supported by JSPS KAKENHI (Grant No. 24K02904), JST CREST (Grant No. JPMJCR21D2), and JST FOREST (Grant No. JPMJFR216I).
@@ B.2.1 Specific Setup for Low-Rank Attention
For example, the original Transformer architecture sets $H=8$ and $n_{H}=64$. Large-scale models often increase the number of heads to be on the order of the hidden embedding dimension [13], as seen in GPT-3 (175B parameters), which sets $H=96$ and $n_{H}=128$.
@@ C.1 Basics
Lemma C.2 (Portmanteau lemma (Lemma 2.2 in [29])).
@@ C.2 Pseudo-Lipschitz Functions
Definition C.1 (Pseudo-Lipschitz functions [6]).
@@ D.1.1 Weak Convergence of the Dot Products
The proof of Proposition D.1 relies on the next lemma, which is an application of Theorem 2 in [5].
@@
$\mathbb{E}(X_{\alpha}^{4})$ $\lesssim\sum_{i_{1},\dots,i_{4}=1}^{r}\sum_{\gamma_{1},\dots,\gamma_{8}=1}^{n}\mathbb{E}\left[W_{\alpha\gamma_{1}}^{i_{1},1}W_{\alpha\gamma_{2}}^{i_{2},1}W_{\alpha\gamma_{3}}^{i_{3},1}W_{\alpha\gamma_{4}}^{i_{4},1}W_{\alpha\gamma_{5}}^{i_{1},2}W_{\alpha\gamma_{6}}^{i_{2},2}W_{\alpha\gamma_{7}}^{i_{3},2}W_{\alpha\gamma_{8}}^{i_{4},2}\right]$
$=\sum_{i=1}^{r}\sum_{\gamma_{1},\dots,\gamma_{8}=1}^{n}\mathbb{E}\left[W_{\alpha\gamma_{1}}^{i,1}W_{\alpha\gamma_{2}}^{i,1}W_{\alpha\gamma_{3}}^{i,1}W_{\alpha\gamma_{4}}^{i,1}W_{\alpha\gamma_{5}}^{i,2}W_{\alpha\gamma_{6}}^{i,2}W_{\alpha\gamma_{7}}^{i,2}W_{\alpha\gamma_{8}}^{i,2}\right]$
$\quad+4\sum_{i_{1}=1}^{r}\sum_{i_{2}\neq i_{1}}\sum_{\gamma_{1},\dots,\gamma_{8}=1}^{n}\mathbb{E}\left[W_{\alpha\gamma_{1}}^{i_{1},1}W_{\alpha\gamma_{2}}^{i_{1},1}W_{\alpha\gamma_{3}}^{i_{1},1}W_{\alpha\gamma_{4}}^{i_{2},1}W_{\alpha\gamma_{5}}^{i_{1},2}W_{\alpha\gamma_{6}}^{i_{1},2}W_{\alpha\gamma_{7}}^{i_{1},2}W_{\alpha\gamma_{8}}^{i_{2},2}\right]$
$\quad+3\sum_{i_{1}=1}^{r}\sum_{i_{2}\neq i_{1}}\sum_{\gamma_{1},\dots,\gamma_{8}=1}^{n}\mathbb{E}\left[W_{\alpha\gamma_{1}}^{i_{1},1}W_{\alpha\gamma_{2}}^{i_{1},1}W_{\alpha\gamma_{3}}^{i_{2},1}W_{\alpha\gamma_{4}}^{i_{2},1}W_{\alpha\gamma_{5}}^{i_{1},2}W_{\alpha\gamma_{6}}^{i_{1},2}W_{\alpha\gamma_{7}}^{i_{2},2}W_{\alpha\gamma_{8}}^{i_{2},2}\right]$
$\quad+6\sum_{i_{1}=1}^{r}\sum_{i_{2}\neq i_{1}}\sum_{i_{3}\notin\{i_{1},i_{2}\}}\sum_{\gamma_{1},\dots,\gamma_{8}=1}^{n}\mathbb{E}\left[W_{\alpha\gamma_{1}}^{i_{1},1}W_{\alpha\gamma_{2}}^{i_{1},1}W_{\alpha\gamma_{3}}^{i_{2},1}W_{\alpha\gamma_{4}}^{i_{3},1}W_{\alpha\gamma_{5}}^{i_{1},2}W_{\alpha\gamma_{6}}^{i_{1},2}W_{\alpha\gamma_{7}}^{i_{2},2}W_{\alpha\gamma_{8}}^{i_{3},2}\right]$
@@
Applying a similar argument, we can also show that $A_{4}\lesssim\sigma^{8}$ and $A_{5}\lesssim\sigma^{8}$ holds. Therefore, we conclude that $\mathbb{E}(X_{\alpha}^{4})\lesssim\sigma^{8}$ holds, which completes the proof. ∎
@@
To upgrade this to convergence in expectation, Facts C.1 and C.2 imply it suffices to show the existance of a constant $\delta>1$ that satisfies

### v3

@@ 5.2 Scalings of the Dot-product Score and Finiteness of Heads
Figure 2: Visualization of the dot-product score $p_{1,1}^{(1)}$ and attention output $y_{1}^{1}$, as defined in Example 3.1, comparing finite-width behavior to their infinite-width limits. (a) Histogram of the empirical distribution of $p_{1,1}^{(1)}$ for $n=256$ alongside the plot of its infinite-width limit distribution $\mathring{p}_{1,1}^{1}$, under two scaling schemes; $1/\sqrt{n}$ and $1/n$. The $1/n$-scaled score collapses to zero in the infinite-width limit, while the $1/\sqrt{n}$-scaled score retains a nondegenerate distribution. (b) Kernel density estimates of the empirical distribution of $y_{1}^{1}$ for $n=256$ (dashed lines) alongside the plot of its infinite-width limit distribution $Z^{y^{1}}$ (solid lines), varying head counts $H\in\{1,256\}$. The black solid line represents the density of the infinite-head limit distribution from hron2020infinite. This demonstrates that our theoretical prediction remains accurate even when $H$ grows, and it approaches the infinite-head limit.
We next investigate two facets of finite-width behavior in Example 3.1, which are the effect of different scaling rules on the variability of the dot-product score and the impact of a finite number of heads on the attention output. For comparison, yang2019tensor1 assumes $1/n$-scaling, and the infinite-head result by hron2020infinite applies only in the limit $n,H\to\infty$.
@@
Figure 2(b) reports histograms of the attention output $y_{1}^{1}$ for $n=256$ with head counts $H=1$ and $H=256$, overlaid with the infinite-width densities of $Z^{y^{1}}$, which are approximated via Monte Carlo sampling. The close agreement between finite-width histograms and their infinite-width limit curves, even when $H$ grows on the same order as $n$, demonstrates the robustness of our infinite-width approximation in the growing number of heads. Furthermore, it is observed that as $H$ increases, our non-Gaussian distribution approaches the Gaussian distribution from hron2020infinite.
@@ 6 Conclusion
Looking forward, we believe our framework can serve as a foundation for future extensions to deep Transformer architectures. We predict that when layers are stacked as in Transformers, not only attention layers but also MLP layers converge to non-Gaussian distributions in the infinite-width limit. The non-Gaussianity of attention layers (and possibly MLP layers) is not merely a theoretical curiosity, but it has profound implications for the learning dynamics of attention-based models. For instance, the presence of higher-order moments associated with such non-Gaussianity could influence signal propagation by preserving feature variability across layers, thereby reducing the risk of rank collapse. Furthermore, the anisotropy induced by non-Gaussianity may lead to more irregular curvature in the optimization landscape, possibly affecting the convergence properties of training dynamics (e.g., through sharper gradients or more prominent saddle regions). Consequently, we argue that a new framework distinct from existing Gaussian-based approaches is essential for analyzing deep architectures. Our work provides an important first step in this direction.
@@ Acknowledgements
We are grateful to the reviewers for their insightful comments. We also thank Yuta Matano for helpful discussions and for pointing out the sub-Gaussianity of the limiting distribution of the attention output. Mana Sakai was supported by RIKEN Junior Research Associate Program. Ryo Karakida was supported by JST FOREST (Grant No. JPMJFR226Q) and JSPS KAKENHI (Grant No. 22H05116). Masaaki Imaizumi was supported by JSPS KAKENHI (Grant No. 24K02904), JST CREST (Grant No. JPMJCR21D2), and JST FOREST (Grant No. JPMJFR216I).
@@ B.2.1 Specific Setup for Low-Rank Attention
For example, the original Transformer architecture sets $H=8$ and $n_{H}=64$. Large-scale models often increase the number of heads to be on the order of the hidden embedding dimension [everett2024scaling], as seen in GPT-3 (175B parameters), which sets $H=96$ and $n_{H}=128$.
@@ C.1 Basics
Lemma C.2 (Portmanteau lemma (Lemma 2.2 in [vaart1998asymptotic])).
@@ C.2 Pseudo-Lipschitz Functions
Definition C.1 (Pseudo-Lipschitz functions [bayati2011dynamics]).
@@ D.1.1 Weak Convergence of the Dot Products
The proof of Proposition D.1 relies on the next lemma, which is an application of Theorem 2 in [blum1958central].
@@
$\mathbb{E}(X_{\alpha}^{4})$ $\lesssim\sum_{i_{1},\dots,i_{4}=1}^{r}\sum_{\gamma_{1},\dots,\gamma_{8}=1}^{n}\mathbb{E}\left[W_{\alpha\gamma_{1}}^{i_{1},1}W_{\alpha\gamma_{2}}^{i_{2},1}W_{\alpha\gamma_{3}}^{i_{3},1}W_{\alpha\gamma_{4}}^{i_{4},1}W_{\alpha\gamma_{5}}^{i_{1},2}W_{\alpha\gamma_{6}}^{i_{2},2}W_{\alpha\gamma_{7}}^{i_{3},2}W_{\alpha\gamma_{8}}^{i_{4},2}\right]\}$
$=\sum_{i=1}^{r}\sum_{\gamma_{1},\dots,\gamma_{8}=1}^{n}\mathbb{E}\left[W_{\alpha\gamma_{1}}^{i,1}W_{\alpha\gamma_{2}}^{i,1}W_{\alpha\gamma_{3}}^{i,1}W_{\alpha\gamma_{4}}^{i,1}W_{\alpha\gamma_{5}}^{i,2}W_{\alpha\gamma_{6}}^{i,2}W_{\alpha\gamma_{7}}^{i,2}W_{\alpha\gamma_{8}}^{i,2}\right]\}$
$\quad+4\sum_{i_{1}=1}^{r}\sum_{i_{2}\neq i_{1}}\sum_{\gamma_{1},\dots,\gamma_{8}=1}^{n}\mathbb{E}\left[W_{\alpha\gamma_{1}}^{i_{1},1}W_{\alpha\gamma_{2}}^{i_{1},1}W_{\alpha\gamma_{3}}^{i_{1},1}W_{\alpha\gamma_{4}}^{i_{2},1}W_{\alpha\gamma_{5}}^{i_{1},2}W_{\alpha\gamma_{6}}^{i_{1},2}W_{\alpha\gamma_{7}}^{i_{1},2}W_{\alpha\gamma_{8}}^{i_{2},2}\right]\}$
$\quad+3\sum_{i_{1}=1}^{r}\sum_{i_{2}\neq i_{1}}\sum_{\gamma_{1},\dots,\gamma_{8}=1}^{n}\mathbb{E}\left[W_{\alpha\gamma_{1}}^{i_{1},1}W_{\alpha\gamma_{2}}^{i_{1},1}W_{\alpha\gamma_{3}}^{i_{2},1}W_{\alpha\gamma_{4}}^{i_{2},1}W_{\alpha\gamma_{5}}^{i_{1},2}W_{\alpha\gamma_{6}}^{i_{1},2}W_{\alpha\gamma_{7}}^{i_{2},2}W_{\alpha\gamma_{8}}^{i_{2},2}\right]\}$
$\quad+6\sum_{i_{1}=1}^{r}\sum_{i_{2}\neq i_{1}}\sum_{i_{3}\notin\{i_{1},i_{2}\}}\sum_{\gamma_{1},\dots,\gamma_{8}=1}^{n}\mathbb{E}\left[W_{\alpha\gamma_{1}}^{i_{1},1}W_{\alpha\gamma_{2}}^{i_{1},1}W_{\alpha\gamma_{3}}^{i_{2},1}W_{\alpha\gamma_{4}}^{i_{3},1}W_{\alpha\gamma_{5}}^{i_{1},2}W_{\alpha\gamma_{6}}^{i_{1},2}W_{\alpha\gamma_{7}}^{i_{2},2}W_{\alpha\gamma_{8}}^{i_{3},2}\right]\}$
@@
Applying a similar argument, we can also show that $A_{4}\lesssim\sigma^{8}$ and $A_{5}\lesssim\sigma^{8}$ holds Therefore, we conclude that $\mathbb{E}(X_{\alpha}^{4})\lesssim\sigma^{8}$ holds, which completes the proof. ∎
@@
To upgrade this to convergence in expectation, Facts C.1 and C.2 imply it suffices to show the existence of a constant $\delta>1$ that satisfies
@@ D.1.3 $S_{2}$ Converges to 0
Appendix E Sub-Gaussianity of the Attention Outputs
In this section, we discuss the sub-Gaussianity of the limiting distribution of the attention outputs from Example 3.1. In this appendix, we provide a proof of the sub-Gaussianity of the single random variable $Z^{y^{i}}$ for any $i\in[s]$, and consider the single-head attention setting (i.e. $H=1$) for simplicity. Specifically, we consider
$Z^{y^{i}}=\sum_{j=1}^{s}\mathrm{SoftMax}_{j}(\mathring{p}_{i,1}^{(1)},\dots,\mathring{p}_{i,s}^{(1)})Z^{\tilde{v}^{1,j}}.$
The same proof strategy also applies to prove the sub-Gaussianity of the vector $(Z^{y^{1}},\dots,Z^{y^{s}})$ with $H$ being any positive integer.
For each $j\in[s]$, define $a_{j}(\mathring{\bm{p}}_{i,1:s}^{(1)})$ by
$a_{j}(\mathring{\bm{p}}_{i,1:s}^{(1)})=\mathrm{SoftMax}_{j}(\mathring{p}_{i,1}^{(1)},\dots,\mathring{p}_{i,s}^{(1)}).$
It follows that
$Z^{y^{i}}|\mathring{\bm{p}}_{i,1:s}^{(1)}\sim\mathcal{N}(0,\bm{a}_{1:s}(\mathring{\bm{p}}_{i,1:s}^{(1)})^{\top}\mathrm{Var}(\bm{Z}^{\tilde{\bm{v}}^{1,1:s}})\bm{a}_{1:s}(\mathring{\bm{p}}_{i,1:s}^{(1)})).$
Since each $a_{j}(\mathring{\bm{p}}_{i,1:s}^{(1)})$ lies in the interval $[0,1]$, we have
$\|\bm{a}_{1:s}(\mathring{\bm{p}}_{i,1:s}^{(1)})\|^{2}\leq\sum_{j=1}^{s}a_{j}(\mathring{\bm{p}}_{i,1:s}^{(1)})=1$
and therefore
$\bm{a}_{1:s}(\mathring{\bm{p}}_{i,1:s}^{(1)})^{\top}\mathrm{Var}(\bm{Z}^{\tilde{\bm{v}}^{1,1:s}})\bm{a}_{1:s}(\mathring{\bm{p}}_{i,1:s}^{(1)})\leq\|\bm{a}_{1:s}(\mathring{\bm{p}}_{i,1:s}^{(1)})\|^{2}\|\mathrm{Var}(\bm{Z}^{\tilde{\bm{v}}^{1,1:s}})\|_{2}\leq\|\mathrm{Var}(\bm{Z}^{\tilde{\bm{v}}^{1,1:s}})\|_{2},$
where $\|\cdot\|_{2}$ is the operator norm. As a result, for any $t\geq 0$, it holds that
$\mathrm{P}(|Z^{y^{i}}|\geq t)$ $=\mathbb{E}\left[\mathrm{P}(|Z^{y^{i}}|\geq t\mid\mathring{\bm{p}}_{i,1:s}^{(1)})\right]$
$\leq 2\mathbb{E}\left[\exp\left(-\frac{t^{2}}{2\bm{a}_{1:s}(\mathring{\bm{p}}_{i,1:s}^{(1)})^{\top}\mathrm{Var}(\bm{Z}^{\tilde{\bm{v}}^{1,1:s}})\bm{a}_{1:s}(\mathring{\bm{p}}_{i,1:s}^{(1)})}\right)\ \Bigr|\ \mathring{\bm{p}}_{i,1:s}^{(1)}\right]$
$\leq 2\exp\left(-\frac{t^{2}}{2\|\mathrm{Var}(\bm{Z}^{\tilde{\bm{v}}^{1,1:s}})\|_{2}}\right)$
by the conditional sub-Gaussianity of $Z^{y^{i}}$ given $\mathring{\bm{p}}_{i,1:s}^{(1)}$.
In this argument, we can also show the sub-Gaussianity of the limiting distribution of attention outputs of the form $y^{i}=\frac{1}{\sqrt{H}}\sum_{a=1}^{H}\sum_{j=1}^{s}\phi_{j}(p_{i,1}^{(a)},\dots,p_{i,s}^{(a)})\tilde{v}^{a,j}$, where the softmax functions are replaced by bounded and pseudo-Lipschitz nonlinearities $\phi_{j}$.